# LiRA: Linguistic Robust Anchoring for Cross-lingual Large Language Models

### v1

@@ 4.1.2 Translation Critic (minimizing $\epsilon_{2}$)
Role for $\epsilon_{2}$. These scores probe adequacy and well-formedness of $y_{k}$ with respect to $x$, serving as a proxy for small semantic divergence between $p(s\!\mid\!x)$ and $p(s\!\mid\!y_{k})$; maximizing their contribution in the policy (see Sec. 4.1.3) drives smaller $\epsilon_{2}$.
4.1.3 Actor–Adaptor
Adaptor. The multilingual encoder and the LLM use different tokenizers ($L_{x}\neq L_{k}$) and embedding spaces. We first apply temporal pooling $P_{r}(\cdot)$ (window/stride $r$, a hyperparameter) and then an MLP to align both sides to a shared dimension $d_{s}$:
$\tilde{\mathbf{u}}=\phi\!\big(W_{\text{mul}}\,P_{r}(\mathrm{Enc}_{\text{multi}}(x))+\mathbf{b}_{\text{mul}}\big),\quad\mathbf{v}_{k}=\psi\!\big(W_{\text{en}}\,P_{r}(\mathrm{Emb}_{\text{LLM}}(y_{k}))+\mathbf{b}_{\text{en}}\big).$ (6)
We define the cross-space similarity
$\mathrm{sim}_{k}\;=\;\cos\!\big(\tilde{\mathbf{u}},\,\mathbf{v}_{k}\big),$ (7)
and add an encoder-alignment term on the chosen candidate $a$:
$\mathcal{L}_{\text{enc}}\;=\;-\,\cos\!\big(\tilde{\mathbf{u}},\,\mathbf{v}_{a}\big).$ (8)
Actor. For each candidate we form the policy feature by concatenating the critic scores with the adaptor similarity:
@@ 4.1.3 Actor–Adaptor
The Actor–Adaptor objective used in training is $\mathcal{L}_{\text{AA}}=\mathcal{L}_{\text{RL}}+\eta\,\mathcal{L}_{\text{enc}}$, which complements the feature-anchoring loss in Sec. 4.1.1
4.1.4 Overall Objective
@@
$\mathcal{L}=\mathcal{L}_{\text{RL}}+\eta\,\mathcal{L}_{\text{enc}}+\lambda\,\mathcal{L}_{\text{anchor}}.$ (12)
Here, $\mathcal{L}_{\text{anchor}}$ reduces the anchoring error $\epsilon_{1}$, while $\mathcal{L}_{\text{trans}}$ favors low-distortion candidates, shrinking $\epsilon_{2}$—jointly tightening the bound in Sec. 3.2.
4.2 LaSR
Given any-language text, a multilingual encoder yields $E_{\text{lr}}$ while a shared English encoder (prompted LLM encoder) yields $E_{\text{en}}$. $E_{\text{lr}}$ denotes the text embedding in the low-resource language, and $E_{\text{en}}$ denotes the embedding of the corresponding English text. We fuse $E_{\text{en}}$ and $E_{\text{lr}}$ into a single $\ell_{2}$-normalized embedding used for both ranking and retrieval. Training is supported by two FIFO buffers: (i) CorrQueue for correlation-based objectives under small batches, and (ii) DocQueue for listwise nDCG with in-language negatives. Both queues are stop-grad for cached entries and are updated in FIFO manner with a maximum size $K$.
Figure 3: An overview of LaSR.
Encoders.
Each branch follows “tokenize $\rightarrow$ encoder $\rightarrow$ pooling”: $E_{\text{en}}=\mathrm{pool}(\mathrm{Enc}_{\text{LLM}}(\text{prompt}(x)))$ and $E_{\text{lr}}=\mathrm{pool}(\mathrm{Enc}_{\text{multi}}(x))$. Pooling is fixed (last-token or mean) and the two streams are concatenated and linearly projected before entering the LLM Transformer. The transformer attends across the two streams and returns a fused hidden vector $\mathbf{z}$, which is then normalized $\hat{\mathbf{z}}=\mathbf{z}/\|\mathbf{z}\|$.
Training steps (Figure 3b).
Each optimization step processes two mini-batches: (1) query batch and (2) document batch. We forward them through the shared encoders and the LLM Transformer to obtain $\{\hat{\mathbf{z}}_{q}\}$ and $\{\hat{\mathbf{z}}_{d}\}$ and compute $s(q,d)$ for the task at hand. Two FIFO buffers are updated in the background:
  • CorrQueue (Rank task). Caches tuples (Pred, Gold) produced on ranking datasets (e.g., STS). At step $t$ we concatenate the current predictions/labels with up to $K$ cached pairs (detached, no gradient) and compute a correlation loss: $\mathcal{L}_{\text{corrQ}}=\alpha\,(1-\mathrm{Pearson})+(1-\alpha)\,(1-\mathrm{Soft\text{-}Spearman})$. See A.4 for details.
  • DocQueue (Retrieval task). Caches (doc_id, language, $\hat{\mathbf{z}}_{d}$) from recent steps for in-language hard-negative mining. Given a query, we form a candidate list (1 positive + mined negatives), compute differentiable ranks and a soft nDCG@k objective $\mathcal{L}_{\text{ndcg}}=1-\mathrm{nDCG}@k$. To avoid mislabeled near negatives, we apply a temperatured down-weighting on very similar non-positives (“safe negatives”), and add two light regularizers (top-1 hinge, mean/variance control) to stabilize training: $\mathcal{L}_{\text{retr}}=\mathcal{L}_{\text{ndcg}}+\lambda_{h}\mathcal{L}_{\text{hinge}}+\lambda_{r}\mathcal{L}_{\text{mv}}$. See A.4 for details.
5 Experiment
5.1 Experimental Details
[…]
Due to potential discrepancies in pretraining data, we ensured fairness by fine-tuning all models used in our experiments (including Qwen3) on each dataset with identical training hyperparameters before evaluation. Interestingly, we observed that fine-tuning Qwen3 brought no significant gains on existing public datasets, except for our newly introduced LazRetrieval dataset. This phenomenon may be attributed to Qwen3 having already seen portions of these public datasets during its pretraining phase.
Table 1: Evaluation on main public retrieval datasets. bold indicates the best result; underlined indicates the second best.
Method MLQARetrieval BelebeleRetrieval STS22 Avg.
@@ 5.1 Experimental Details
LiRA ours 81.66 87.03 75.00 81.23
Method Bd Id My Pk Th Ph Vn Avg.
Sentence-T5-XXL 2021 34.11 71.77 49.19 27.84 23.58 84.20 28.61 44.56
GTR-XXL 2021 34.85 75.92 49.36 30.39 22.94 84.94 46.15 48.17
SimCSE 2021 31.76 66.71 43.27 27.68 32.32 74.00 44.16 44.88
Contriever 2022 39.95 74.95 48.90 35.71 15.43 83.74 64.75 51.00
GTE-large 2023 39.36 77.59 51.88 36.76 17.21 87.65 65.34 52.61
BGE-en-v1.5 2023 41.06 78.78 53.28 37.52 18.35 88.18 68.72 54.09
E5-large-v2 2024 41.04 78.78 53.77 37.22 17.63 87.24 61.31 52.84
E5-Mistral-7B 2024 48.27 75.43 71.01 53.62 61.75 83.18 65.44 64.51
Qwen3-E-8B 2025 65.97 79.18 80.56 65.39 82.48 86.59 77.31 76.78
LiRA ours 66.30 78.53 81.54 68.53 83.12 87.44 78.48 77.71
5.2 Datasets
We use standard public datasets: BelebeleRetrievalBandarkar et al. (2024), MLQARetrievalet al. (2025), STS22et al. (2025), MGSMShi et al. (2022), and X-CSQALin et al. (2021). The first two are retrieval benchmarks, STS22 evaluates sentence-level correlation, and MGSM/X-CSQA assess mathematical reasoning and reading comprehension, respectively. We additionally release a de-identified e-commerce retrieval dataset, LazRetrieval, and a larger companion set, LazRetrieval-mega. Both cover seven Southeast-Asian languages: Vietnamese (Vi), Thai (Th), Indonesian (Id), Malay (Ms), Urdu (Ur), Bengali (Bn), and Filipino/Tagalog (Ph). LazRetrieval contains $\mathbf{10\,k}$ examples per language; LazRetrieval-mega contains $\mathbf{1{,}000\,k}$ examples per language and is intended for pretraining/supporting large-scale adaptation. Unless otherwise noted, our experiments use LazRetrieval. Since MGSM and X-CSQA have no training split in our setup, we evaluate in a zero-shot fashion: the Arca is trained on BelebeleRetrieval, while the lightweight LaSR head is left untrained for these tasks. Detailed information about the datasets can be found in the  A.2.
@@ 5.2 Datasets
Method Bn Th Sw Ja Zh De Fr Ru Es En Avg.
MonoReason 2024 6.8 7.2 6.8 36.4 38.4 55.2 54.4 52.0 57.2 68.8 38.3
MultiReason-Lora 2022 29.6 35.2 28.0 52.0 54.8 59.6 58.4 62.4 59.6 64.8 50.4
MultiReason-SFT 2024 33.2 40.0 42.0 42.0 42.0 45.2 44.8 45.2 48.0 52.0 43.4
QAlign 2024 39.6 40.4 44.0 44.0 48.4 54.8 56.8 52.4 59.6 68.0 49.6
LangBridge 2024 42.8 50.4 43.2 40.0 45.2 56.4 50.8 52.4 58.0 63.2 50.2
Translate-En 2023 48.4 37.6 37.6 49.2 46.8 60.4 56.4 47.6 59.6 65.5 50.6
MindMerger-Hard 2024 46.0 36.0 48.4 52.4 54.4 60.4 56.0 60.4 62.0 71.2 54.7
MindMerger-Soft 2024 50.4 52.8 57.2 54.4 53.6 61.2 57.6 60.8 58.4 66.8 57.3
Qwen3-14B 2025 87.2 90.0 78.0 86.4 88.8 90.0 88.4 92.8 94.0 95.2 89.1
LiRA Ours 88.4 91.2 78.0 86.8 89.6 89.6 88.4 92.8 92.4 95.2 89.2
@@
Qwen3-14B 2025 34.8 50.4 51.6 60.6 60.9 58.3 61.2 65.0 63.6 62.9 66.5 65.8 67.5 65.7 68.2 81.7 61.5
LiRA Ours 35.3 51.1 52.1 61.2 61.6 58.7 61.9 65.7 64.0 63.0 66.1 66.2 67.6 66.2 68.9 82.5 62.0
@@ 5.3 Results Analysis
Retrieval & sentence ranking.
On public benchmarks (Table 1), LiRA consistently improves the base model (Qwen3-E-8B) on all three metrics: MLQARetrieval $81.66$ vs. $81.13$ (+0.53), BelebeleRetrieval $87.03$ vs. $85.94$ (+1.09), and STS22 $75.00$ vs. $71.64$ (+3.36), yielding a higher macro average 81.23 (+1.66). On our new LazRetrieval-70K (Table 2), LiRA also improves the average from $76.78$ to 77.71 (+0.93). The gains are particularly pronounced on relatively low-resource locales (e.g., pk +3.14, vn +1.17, my +0.98), suggesting that anchoring $g(x)$ to the English space and the LaSR head together reduce translation/representation noise. We also observe better rank-correlation (STS22), matching our design of queue-augmented CorrQ and listwise soft-nDCG.
Mathematic.
On MGSM(Table 3), LiRA brings a small but consistent gain over Qwen3-14B: macro avg 89.2 vs. $89.1$ (+0.1). Per-language analysis shows improvements or ties on $9/11$ languages (e.g., Bn/Th/Zh), while a few high-resource languages remain similar (De, Es). This indicates that the concatenated representation $[g(x);h(y)]$ improves reasoning robustness without hurting English performance.
Comprehension.
On X-CSQA(Table 4), LiRA outperforms Qwen3-14B on 15/16 languages and raises the macro average from $61.5$ to 62.0 (+0.5). Improvements concentrate on lower-resource or typologically distant languages (Ur/Hi/Ar/Vi/Zh), consistent with our motivation that concatenating $g(x)$ and $h(y)$ adds complementary information and mitigates information bottlenecks.

### v2

@@ 4.1.2 Translation Critic (minimizing $\epsilon_{2}$)
Role for $\epsilon_{2}$. These scores probe adequacy and well-formedness of $y_{k}$ with respect to $x$, serving as a proxy for small semantic divergence between $p(s\!\mid\!x)$ and $p(s\!\mid\!y_{k})$; maximizing their contribution in the policy drives smaller $\epsilon_{2}$.
4.1.3 Overall Objective
For each candidate we form the policy feature by concatenating the critic scores with the adaptor similarity:
@@ 4.1.3 Overall Objective
$\mathcal{L}=\mathcal{L}_{\text{RL}}+\eta\,\mathcal{L}_{\text{anchor}}.$ (9)
Here, $\mathcal{L}_{\text{anchor}}$ reduces the anchoring error $\epsilon_{1}$, while $\mathcal{L}_{\text{RL}}$ favors low-distortion candidates, shrinking $\epsilon_{2}$—jointly tightening the bound in Sec. 3.2.
Method Params Bd Id My Pk Th Ph Vn Avg.
Sentence-T5-XXL 2021 4.8B 34.11 71.77 49.19 27.84 23.58 84.20 28.61 44.56
GTR-XXL 2021 4.8B 34.85 75.92 49.36 30.39 22.94 84.94 46.15 48.17
SimCSE 2021 330M 31.76 66.71 43.27 27.68 32.32 74.00 44.16 44.88
Contriever 2022 110M 39.95 74.95 48.90 35.71 15.43 83.74 64.75 51.00
GTE-large 2023 335M 39.36 77.59 51.88 36.76 17.21 87.65 65.34 52.61
BGE-en-v1.5 2023 335M 41.06 78.78 53.28 37.52 18.35 88.18 68.72 54.09
E5-large-v2 2024 335M 41.04 78.78 53.77 37.22 17.63 87.24 61.31 52.84
E5-Mistral-7B 2024 7.24B 48.27 75.43 71.01 53.62 61.75 83.18 65.44 64.51
Qwen3-E-0.6B 2025 0.6B 38.36 63.95 62.37 40.73 55.28 74.38 59.46 56.31
With LiRA-Large ours 8.5B 48.60 74.43 71.26 49.84 66.39 83.90 70.67 66.44
Qwen3-E-4B 2025 4B 49.81 68.92 71.45 49.30 73.39 78.47 68.31 65.66
With LiRA-Large ours 11.9B 56.24 75.29 77.03 55.69 77.72 84.81 74.92 71.67
Qwen3-E-8B 2025 8B 50.59 76.01 73.36 50.37 69.67 84.78 71.56 68.05
With LiRA-Base ours 14.4B 52.91 76.59 75.20 52.30 71.11 85.92 73.23 69.61
With LiRA-Large ours 15.9B 57.70 77.33 77.93 56.67 78.52 86.03 75.86 72.86
With LiRA-Max ours 103B 66.30 78.53 81.54 68.53 83.12 87.44 78.48 77.71
Table 2: Evaluation on main public retrieval datasets, and the effect of adding LiRA-Max to representative embedding backbones. bold indicates the best; underlined indicates the second best. Column abbreviations: MLQA-R = MLQARetrieval, Bele-R = BelebeleRetrieval.
Method MLQA-R Bele-R STS22 Avg.
@@
$\hookrightarrow$ with LiRA 82.01 87.03 75.00 81.35
4.2 LaSR
Given any-language text, a multilingual encoder yields $E_{\text{lr}}$ while a shared English encoder (prompted LLM encoder) yields $E_{\text{en}}$. $E_{\text{lr}}$ denotes the text embedding in the low-resource language, and $E_{\text{en}}$ denotes the embedding of the corresponding English text. We fuse $E_{\text{en}}$ and $E_{\text{lr}}$ into a single $\ell_{2}$-normalized embedding used for all ranking, retrieval and reasoning tasks. Training is supported by two FIFO buffers: (i) CorrQueue for correlation-based objectives under small batches, and (ii) DocQueue for listwise nDCG with in-language negatives. Both queues are stop-grad for cached entries and are updated in FIFO manner with a maximum size $K$.
Encoders. Each branch follows “tokenize $\rightarrow$ encoder $\rightarrow$ pooling”: $E_{\text{en}}$ and $E_{\text{lr}}$, as shwon in Eq. (4). Pooling is fixed and the two streams are concatenated and linearly projected before entering the LLM Transformer. The transformer attends across the two streams and returns a fused hidden vector $\mathbf{z}$, which is then normalized $\hat{\mathbf{z}}=\mathbf{z}/\|\mathbf{z}\|$.
Training steps (Figure 3b). Each optimization step processes two mini-batches: (1) query batch and (2) document batch. We forward them through the shared encoders and the LLM Transformer to obtain $\{\hat{\mathbf{z}}_{q}\}$ and $\{\hat{\mathbf{z}}_{d}\}$ and compute $s(q,d)$ for the task at hand. Two FIFO buffers are updated in the background:
  • CorrQueue (Rank task). Caches tuples (Pred, Gold) produced on ranking datasets (e.g., STS). At step $t$ we concatenate the current predictions/labels with up to $K$ cached pairs (detached, no gradient) and compute a correlation loss: $\mathcal{L}_{\text{corrQ}}=\alpha\,(1-\mathrm{Pearson})+(1-\alpha)\,(1-\mathrm{Soft\text{-}Spearman})$. See Appendix D.4 for details.
  • DocQueue (Retrieval task). Caches (doc_id, language, $\hat{\mathbf{z}}_{d}$) from recent steps for in-language hard-negative mining. Given a query, we form a candidate list (1 positive + mined negatives), compute differentiable ranks and a soft nDCG@k objective $\mathcal{L}_{\text{ndcg}}=1-\mathrm{nDCG}@k$. To avoid mislabeled near negatives, we apply a temperatured down-weighting on very similar non-positives (“safe negatives”), and add two light regularizers (top-1 hinge, mean/variance control) to stabilize training: $\mathcal{L}_{\text{retr}}=\mathcal{L}_{\text{ndcg}}+\lambda_{h}\mathcal{L}_{\text{hinge}}+\lambda_{r}\mathcal{L}_{\text{mv}}$. See Appendix D.4 for details.
@@ 4.2 LaSR
Method Params Year Bn Th Sw Ja Zh De Fr Ru Es En Avg.
MonoReason 7B 2024 6.8 7.2 6.8 36.4 38.4 55.2 54.4 52.0 57.2 68.8 38.3
MultiReason-Lora 7B 2022 29.6 35.2 28.0 52.0 54.8 59.6 58.4 62.4 59.6 64.8 50.4
MultiReason-SFT 7B 2024 33.2 40.0 42.0 42.0 42.0 45.2 44.8 45.2 48.0 52.0 43.4
QAlign 7B 2024 39.6 40.4 44.0 44.0 48.4 54.8 56.8 52.4 59.6 68.0 49.6
LangBridge 7B 2024 42.8 50.4 43.2 40.0 45.2 56.4 50.8 52.4 58.0 63.2 50.2
Translate-En 3.3B 2023 48.4 37.6 37.6 49.2 46.8 60.4 56.4 47.6 59.6 65.5 50.6
MindMerger-Hard 10.7B 2024 46.0 36.0 48.4 52.4 54.4 60.4 56.0 60.4 62.0 71.2 54.7
MindMerger-Soft 10.7B 2024 50.4 52.8 57.2 54.4 53.6 61.2 57.6 60.8 58.4 66.8 57.3
Qwen3-8B 8B 2025 66.4 69.3 59.1 64.5 67.9 71.0 69.3 72.5 75.1 78.9 69.4
With LiRA-Large 15.9B Ours 69.6 72.5 61.9 69.1 70.3 69.3 69.8 73.9 75.0 79.2 71.1
@@
Qwen3-8B 2025 35.7 51.6 52.8 60.9 63.0 59.3 62.5 66.6 64.7 64.1 67.6 66.9 68.2 69.8 70.1 82.8 62.9
With LiRA-Large Ours 40.8 52.7 55.6 63.9 65.0 61.3 64.2 68.3 67.9 66.3 69.7 70.8 72.0 70.7 74.6 84.2 65.5
4.3 Motivation and Discussion.
A key challenge in low-resource multilingual reasoning is the representation imbalance across languages: modern LLMs typically acquire substantially stronger reasoning capability in high-resource languages (notably English), where reasoning-oriented supervision is abundant, while comparable signals are scarce in many low-resource languages. Under this data constraint, our goal is to make the most of the strong reasoning capabilities that LLMs have acquired in high-resource languages and transfer them to underrepresented languages in a robust and analyzable manner.
LiRA addresses the imbalance by explicitly modeling and shrinking two ubiquitous sources of cross-lingual shift that degrade downstream retrieval and reasoning: (i) semantic drift caused by translation noise and linguistic variation, and (ii) representation mapping error between multilingual and English embedding spaces. Arca reduces both shifts via anchor-based alignment and critic–actor training, which aims to mitigate translation-induced noise rather than propagate it. Importantly, LiRA is plug-and-play: it can be attached to different pretrained backbones with lightweight fine-tuning, and it does not assume ideal translations in practice. On top of Arca, LaSR performs a two-way coupling between multilingual and English representations with queue-based objectives, balancing cross-lingual alignment and task-driven discrimination so that English-centric reasoning signals can be leveraged without destabilizing multilingual embeddings.
5 Experiment
5.1 Experimental Details
[…]
Due to potential discrepancies in pretraining data, we ensured fairness by fine-tuning all models used in our experiments on each dataset with identical training hyperparameters before evaluation. Interestingly, we observed that fine-tuning Qwen3 brought no gains on existing public datasets. This phenomenon may be attributed to Qwen3 having already seen portions of these public datasets during its pretraining phase. For the critic, the weights are set to $\alpha,\beta,\gamma,\delta=0.4,0.4,0.3,1.0$.
We instantiate LiRA at three parameter scales. LiRA-Base uses three lightweight MT models—OPUS-MT, m2m100-418M, and nllb-200-600M—together with Qwen3-1.7B, which also serves as the critic. LiRA-Large upgrades the translator set to Qwen3-1.7B, DeepSeek-R1-Distill-Qwen-1.5B, and Llama3.2-1B-Instruct, with Qwen3-1.7B again acting as the critic model. Finally, LiRA-Max employs three large LLMs—Qwen3-32B, DeepSeek-R1-Distill-Qwen-32B, and Gemma-2-27B—where Qwen3-32B is used as the critic. See D.1 for more detailed information.
[…]
5.2 Datasets
We use standard public datasets: BelebeleRetrieval (Bandarkar et al., 2024), MLQARetrieval (Enevoldsen et al., 2025), STS22 (Enevoldsen et al., 2025), MGSM (Shi et al., 2022), and X-CSQA (Lin et al., 2021). The first two are retrieval benchmarks, STS22 evaluates sentence-level correlation, and MGSM/X-CSQA assess mathematical reasoning and reading comprehension, respectively. As we have observed above that contamination in cross-lingual training data may inflate the reported performance, we additionally introduce a new real-world dataset to enable a fair comparison. We release a de-identified e-commerce retrieval dataset, LazRetrieval, and a larger companion set, LazRetrieval-mega for further research. Both cover seven Southeast-Asian languages: Vietnamese (Vi), Thai (Th), Indonesian (Id), Malay (Ms), Urdu (Ur), Bengali (Bn), and Filipino/Tagalog (Ph). LazRetrieval contains $\mathbf{10\,k}$ examples per language; LazRetrieval-mega contains $\mathbf{1{,}000\,k}$ examples per language and is intended for pretraining/supporting large-scale adaptation. Unless otherwise noted, our experiments use LazRetrieval. Since MGSM and X-CSQA have no training split in our setup, we evaluate in a zero-shot fashion: the Arca is trained on BelebeleRetrieval, while the lightweight LaSR head is left untrained for these tasks. Detailed information about the datasets can be found in  B.
@@ 5.3 Results Analysis
Retrieval & sentence ranking. On public benchmarks (Table 2), LiRA consistently improves the base model (Qwen3-E-8B) on all three metrics: MLQARetrieval $82.01$ vs. $81.13$ (+0.88), BelebeleRetrieval $87.03$ vs. $85.94$ (+1.09), and STS22 $75.00$ vs. $71.64$ (+3.36), yielding a higher macro average 81.35 (+1.78). On our new LazRetrieval-70K (Table 1), Qwen3-Embedding with LiRA-Large also improves the average from $68.05$ to 72.86 (+4.81). The gains are particularly pronounced on relatively low-resource locales (e.g., pk +6.40, vn +4.30, my +4.57), suggesting that anchoring $g(x)$ to the English space and the LaSR head together reduce translation/representation noise. We also observe better rank-correlation, matching our design of queue-augmented CorrQ and listwise soft-nDCG.
Scaling behaviour on LazRetrieval. On LazRetrieval, we observe a consistent scaling trend across Qwen3 embedding backbones of different sizes (Tab. 1). For the smallest backbone (Qwen3-E-0.6B), attaching LiRA-Large yields substantial gains on low-resource languages such as Bd and Pk (e.g., +10.24 and +9.11 absolute points, respectively), indicating that LiRA can effectively compensate for limited model capacity. A similar pattern holds for Qwen3-E-4B, where LiRA-Large improves Bd from 49.81 to 56.24 and Pk from 49.30 to 55.69. For the full 8B backbone, LiRA forms a smooth performance–capacity curve: the average score increases from 68.05 to 69.61 with LiRA-Base, 72.86 with LiRA-Large, and 77.71 with LiRA-Max, accompanied by consistent improvements on all seven LazRetrieval languages. These results demonstrate that LiRA provides monotonic benefits across parameter scales and is particularly effective in enhancing smaller cross-lingual encoders.
Mathematic. On MGSM (Table 3), LiRA brings a small but consistent gain over Qwen3-8B: macro avg 69.4 vs. $71.1$ (+1.7). Per-language analysis shows improvements or matches on $9/11$ languages (e.g., Bn/Th/Zh), while performance on several high-resource languages remains comparable (De, Es). This indicates that the concatenated representation $[g(x);h(y)]$ improves reasoning robustness without compromising performance on high-resource languages.
Comprehension. On X-CSQA (Table 4), LiRA outperforms Qwen3-8B on 15/16 languages and raises the macro average from $62.9$ to 65.5 (+2.6). Improvements concentrate on lower-resource or typologically distant languages (Ur/Hi/Ar/Vi/Zh), consistent with our motivation that concatenating $g(x)$ and $h(y)$ adds complementary information and mitigates information bottlenecks.
Cross-backbone robustness. Table 2 also evaluates LiRA as a pluggable module on three representative encoders. Across MLQA Retrieval, BelebeleRetrieval, and STS22, LiRA consistently improves over the corresponding backbones. The averaged gains over the three tasks are positive for all backbones tested, suggesting that the effect is not tied to a single encoder family.